OpenStudy (omgthatgirl):

Which number completes the square? X^2-20x+? A. -100 B. -10 C. 10 <--- my answer D. 100

OpenStudy (perl):

$$ \Large{ x^2 + bx + \color{red}{(\frac b2)^2} = ( x + \frac b 2 )^2 \\ \therefore \\ x^2 + (-20)\cdot x +\color{red}{(\frac{-20}{2}) ^2} = ( x + \frac{-20}{2})^2 } $$

OpenStudy (omgthatgirl):

Ok so its -10

OpenStudy (perl):

don't forget to square it :)

OpenStudy (omgthatgirl):

100?

OpenStudy (perl):

yes